# Sampling from the Mean-Field Stationary Distribution

### v2

@@ Abstract
We study the complexity of sampling from the stationary distribution of a mean-field SDE, or equivalently, the complexity of minimizing a functional over the space of probability measures which includes an interaction term. Our main insight is to decouple the two key aspects of this problem: (1) approximation of the mean-field SDE via a finite-particle system, via uniform-in-time propagation of chaos, and (2) sampling from the finite-particle stationary distribution, via standard log-concave samplers. Our approach is conceptually simpler and its flexibility allows for incorporating the state-of-the-art for both algorithms and theory. This leads to improved guarantees in numerous settings, including better guarantees for optimizing certain two-layer neural networks in the mean-field regime.
@@ 1 Introduction
The minimization of energy functionals $\mathcal{E}$ over the Wasserstein space $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ of probability measures has attracted substantial research activity in recent years, encompassing numerous application domains, including distributionally robust optimization [Kuh+19, YKW22], sampling [JKO98, Wib18, Che23], and variational inference [LW16, Lam+22, Dia+23, JCP23, Lac23, YY23].
A canonical example of such a functional is $\mathcal{E}(\mu)=\int V\,\mathrm{d}\mu+\int\log\mu\,\mathrm{d}\mu$, where $V:\mathbb{R}^{d}\to\mathbb{R}$ is called the potential. Up to an additive constant, which is irrelevant for the optimization, this energy functional equals the KL divergence $\operatorname{\mathsf{KL}}(\mu\mathbin{\|}\pi)$ with respect to the density $\pi\propto\exp(-V)$, and the celebrated result of [JKO98] identifies the Wasserstein gradient flow of $\mathcal{E}$ with the Langevin diffusion. This link has inspired a well-developed theory for log-concave sampling, with applications to Bayesian inference and randomized algorithms; see [Che23] for an exposition.
The energy functional above contains two terms, corresponding to two of the fundamental examples of functionals considered in Villani’s well-known treatise on optimal transport [Vil03]. Namely, they are the “potential energy” and the entropy, the latter being a special case of the “internal energy.” However, Villani identifies a third fundamental functional—the “interaction energy”—with the pairwise form given by
@@
where $\mathcal{F}:\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})\to\mathbb{R}$ is a known functional. The minimization of the energy ($\mathsf{gE}$) has recently been of interest due to its role in analysing neural network training dynamics in the mean-field regime, including with [SNW22] and without [CB18, MMN18] entropic regularization, as well as with Fisher regularization [Cla+23].
@@
However, here too there is a well-developed mathematical theory which suggests a principled algorithmic approach. Just as the Wasserstein gradient flow of ($\mathsf{pE}$) in the case when $W=0$ can be identified with the Langevin diffusion, the Wasserstein gradient flow of ($\mathsf{pE}$) in the case when $W\neq 0$ corresponds to a (pairwise) McKean–Vlasov SDE, i.e., an SDE whose coefficients depend on the marginal law of the process, given as
@@
where $\{B^{i}:i\in[N]\}$ is a collection of independent Brownian motions. Moreover, the error from approximating the mean-field limit via this finite particle system has been studied in the literature on propagation of chaos [Szn91]. Similarly, the Wasserstein gradient flow for ($\mathsf{gE}$) corresponds to the mean-field Langevin dynamics and admits an analogous particle approximation.
The bounds for propagation of chaos have been refined over time, with [LL23] recently establishing a tight error dependence $\mathcal{O}(\nicefrac{{1}}{{N}})$ on the total number of particles $N$. These bounds, however, do not translate immediately into algorithmic guarantees. Existing sampling analyses study the propagation of chaos and discretization as a single entangled problem, which thus far have only been able to use weaker ${\mathcal{O}}(\sqrt{\nicefrac{{1}}{{N}}})$ rates for the former. Furthermore, there has been recent interest in using more sophisticated particle-based algorithms, e.g., “non-linear” Hamiltonian Monte Carlo [BS23] and the mean-field underdamped Langevin dynamics [FW23] to reduce the discretization error. Currently, this requires repeatedly carrying out the propagation of chaos and time discretization analyses from the ground up for each instance.
This motivates us to pose the following questions: (1) Can we incorporate improvements in the propagation of chaos literature, such as the $\mathcal{O}(\nicefrac{{1}}{{N}})$ error dependence shown in [LL23], to improve existing theoretical guarantees? (2) Can we leverage recent advances in the theory of log-concave sampling to design better algorithms?
Our main proposal in this work is to decouple the error into two terms, representing the propagation of chaos and discretization errors respectively. This simple and modular approach immediately allows us to answer both questions in the affirmative. Namely, we show how to combine established propagation of chaos bounds in various settings [[, including the sharp rate of]]lacker2023sharp with a large class of sophisticated off-the-shelf log-concave samplers, such as interacting versions of the randomized midpoint discretization of the underdamped Langevin dynamics [SL19, HBE20], Metropolis-adjusted algorithms [Che+21, WSC22, AC23], and the proximal sampler [LST21, Che+22, FYC23]. Our framework yields improvements upon prior state-of-the-art, such as [BS23, FW23], and provides a clear path for future ones.
@@ 1.1 Contributions and Organization
We provide three propagation of chaos results which hold in the $\mathcal{W}_{2}$, $\sqrt{\operatorname{\mathsf{KL}}}$, and $\sqrt{\operatorname{\mathsf{FI}}}$ “metrics”; the rates reflect the distance of the $k$-particle marginal of the finite-particle system from $\pi^{\otimes k}$: (1) In the setting of ($\mathsf{pE}$), under strong displacement convexity, we obtain a $\mathcal{O}(\sqrt{\nicefrac{{k}}{{N}}})$ rate by adapting techniques from [Szn91, Mal01]; (2) without assuming displacement convexity, but assuming a weaker interaction, we obtain the sharp rate of $\widetilde{\mathcal{O}}(\nicefrac{{k}}{{N}})$ following [LL23]; (3) finally, in the general setting of ($\mathsf{gE}$), and assuming $\mathcal{F}$ is convex along linear interpolations, we obtain a $\mathcal{O}(\sqrt{\nicefrac{{k}}{{N}}})$ rate following [CRW22].
@@ 1.2 Related Work
The McKean–Vlasov SDE was first formulated in the works [McK66, Fun84, Mél96], with origins dating to much earlier [Bol72]. It has applications in many domains, from fluid dynamics [Vil02] to game theory [LL07, CD18]; see [CD22, CD22a] for a comprehensive survey. The kinetic version of this equation is known as the Boltzmann equation, and propagation of chaos has similarly been studied under a variety of assumptions [BGM10, Mon17, GM21, GLM22]. One prominent application within machine learning is the study of infinitely wide two-layer neural networks in the mean-field regime (see §4.2).
@@
The original propagation of chaos arguments of [Szn91] were first made uniform in time in [Mal01, Mal03] in both entropy and $\mathcal{W}_{2}$. The aforementioned works all achieve an error of order $\widetilde{\mathcal{O}}(\sqrt{\nicefrac{{k}}{{N}}})$, and require a strong convexity assumption on $V$ and $W$. These were later adapted for non-smooth potentials [JW17, JW18, BJW23]. Finally, [CRW22] obtained an entropic propagation of chaos bound under a higher-order smoothness assumption. See [CD22] for a more complete bibliography.
The breakthrough result of [Lac21] obtained the sharp bound of $\widetilde{\mathcal{O}}(\nicefrac{{k}}{{N}})$ when the interaction is sufficiently weak, and this bound was made uniform in time in [LL23]. Their approach differs significantly from previous proofs by considering a local analysis based on the recursive BBGKY hierarchy. These results have been extended to other divergences, e.g., the $\chi^{2}$ divergence, but without a uniform-in-time guarantee [HR23].
The question of sampling from minimizers of ($\mathsf{pE}$) was first studied in [Tal96, BT97, AK02]. These works focused on the Euler–Maruyama discretization of the finite-particle system ($\mathsf{pMV}_{N}$), under $L^{\infty}$-boundedness of the gradients. Subsequently, the convergence of the Euler–Maruyama scheme has been studied in many works, including but not limited to [BH22, RES22, Li+23]. Finally, [BS23] considered a non-linear version of Hamiltonian Monte Carlo; we give a detailed comparison with their work in §4.
@@
The mean-field (underdamped) Langevin algorithm for minimizing ($\mathsf{gE}$) was proposed and studied in [CRW22, Che+23]. Under alternative assumptions (see §3.1.2), they established propagation of chaos with a $\mathcal{O}(\sqrt{\nicefrac{{k}}{{N}}})$ rate, for both the overdamped and the underdamped finite-particle approximations. Recent works from the machine learning community [NWS22, SNW22, FW23, SWN23] studied the application of these algorithms for optimizing two-layer neural networks and obtained sampling guarantees. We provide a detailed comparison with their works in §4.2.
@@ 2 Background and Notation
When $\log(1/\pi)$ is $\alpha$-uniformly convex for $\alpha>0$, it follows from the Bakry–Émery condition that $\pi$ satisfies (LSI) with constant $C_{\mathsf{LSI}}\leq 1/\alpha$ [BGL14, Proposition 5.7.1].
@@
Lastly, we recall that the celebrated Otto calculus interprets the space $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, equipped with the $\mathcal{W}_{2}$ metric, as a formal Riemannian manifold [Ott01]. In particular, the Wasserstein gradient of a functional $\mathcal{L}:\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})\to\mathbb{R}\cup\{\infty\}$ is given as $\nabla_{\mathcal{W}_{2}}\mathcal{L}=\nabla\delta\mathcal{L}$. Here, $\delta\mathcal{L}$ is the first variation defined as follows: for all $\nu_{0},\nu_{1}\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, $\delta\mathcal{L}(\nu_{0}):\mathbb{R}^{d}\to\mathbb{R}$ satisfies
@@
The first variation is defined up to an additive constant, but the Wasserstein gradient is unambiguous. See [AGS08] for a rigorous development. As a shorthand, we will write $\delta\mathcal{L}(\nu_{0},x)\coloneqq\delta\mathcal{L}(\nu_{0})(x)$ and similarly $\nabla_{\mathcal{W}_{2}}\mathcal{L}(\nu_{0},x)\coloneqq\nabla_{\mathcal{W}_{2}}\mathcal{L}(\nu_{0})(x)$.
@@ 2.1.1 The Pairwise McKean–Vlasov Setting
In the formalism introduced in the previous section, we note that ($\mathsf{pMV}$) can be interpreted as Wasserstein gradient flow for ($\mathsf{pE}$). In this paper, we refer to ($\mathsf{pMV}$) as the pairwise McKean–Vlasov process. As noted in the introduction, it has the stationary distribution (1.1) which minimizes ($\mathsf{pE}$). Recall also that the equation ($\mathsf{pMV}$) is the mean-field limit of the finite-particle system ($\mathsf{pMV}_{N}$). This $N$-particle system has the following stationary distribution: for $x^{1:N}=[x^{1},\dotsc,x^{N}]\in\mathbb{R}^{d\times N}$,
@@ 2.1.2 The General McKean–Vlasov Setting
More generally, we consider the functional ($\mathsf{gE}$) where $\mathcal{F}$ is of the form $\mathcal{F}(\mu)=\mathcal{F}_{0}(\mu)+\frac{\lambda}{2}\int\lVert\cdot\rVert^{2}\,\mathrm{d}\mu$ with $\lambda\geq 0$. The second term acts as regularization and is common in the literature [FW23, SWN23]. We can describe its Wasserstein gradient flow as the marginal law of a particle trajectory satisfying the following SDE, which we call the general McKean–Vlasov equation:
@@
and $\rho_{x^{1:N}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}$ is the empirical measure of the particle system. The stationary distribution for ($\mathsf{gMV}_{N}$) is given as follows [CRW22, (2.16)]: for $x^{1:N}=[x^{1},\dotsc,x^{N}]\in\mathbb{R}^{d\times N}$,

### v4

@@ Abstract
We study the complexity of sampling from the stationary distribution of a mean-field SDE, or equivalently, the complexity of minimizing a functional over the space of probability measures which includes an interaction term. Our main insight is to decouple the two key aspects of this problem: (1) approximation of the mean-field SDE via a finite-particle system, via uniform-in-time propagation of chaos, and (2) sampling from the finite-particle stationary distribution, via standard log-concave samplers. Our approach is conceptually simpler and its flexibility allows for incorporating the state-of-the-art for both algorithms and theory. This leads to improved guarantees in numerous settings, including better guarantees for optimizing certain two-layer neural networks in the mean-field regime. A key technical contribution is to establish a new uniform-in-$N$ log-Sobolev inequality for the stationary distribution of the mean-field Langevin dynamics.
@@ 1 Introduction
The minimization of energy functionals $\mathcal{E}$ over the Wasserstein space $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ of probability measures has attracted substantial research activity in recent years, encompassing numerous application domains, including distributionally robust optimization [Kuhn+19DRO, YueKuhWie22OptWass], sampling [jordan1998variational, Wib18SamplingOpt, ChewiBook], and variational inference [LiuWan16SVGD, Lam+22GVI, Diao+23FBGVI, JiaChePoo23MFVI, Lac23MFVI, YaoYan23MFVI].
A canonical example of such a functional is $\mathcal{E}(\mu)=\int V\,\mathrm{d}\mu+\int\log\mu\,\mathrm{d}\mu$, where $V:\mathbb{R}^{d}\to\mathbb{R}$ is called the potential. Up to an additive constant, which is irrelevant for the optimization, this energy functional equals the KL divergence $\operatorname{\mathsf{KL}}(\mu\mathbin{\|}\pi)$ with respect to the density $\pi\propto\exp(-V)$, and the celebrated result of [jordan1998variational] identifies the Wasserstein gradient flow of $\mathcal{E}$ with the Langevin diffusion. This link has inspired a well-developed theory for log-concave sampling, with applications to Bayesian inference and randomized algorithms; see [ChewiBook] for an exposition.
The energy functional above contains two terms, corresponding to two of the fundamental examples of functionals considered in Villani’s well-known treatise on optimal transport [villani2003topics]. Namely, they are the “potential energy” and the entropy, the latter being a special case of the “internal energy.” However, Villani identifies a third fundamental functional—the “interaction energy”—with the pairwise form given by
@@
where $\mathcal{F}:\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})\to\mathbb{R}$ is a known functional. The minimization of the energy ($\mathsf{gE}$) has recently been of interest due to its role in analysing neural network training dynamics in the mean-field regime, including with [suzuki2022uniform] and without [chizat2018global, mei2018mean] entropic regularization, as well as with Fisher regularization [Cla+23MeanFieldFI].
@@
However, here too there is a well-developed mathematical theory which suggests a principled algorithmic approach. Just as the Wasserstein gradient flow of ($\mathsf{pE}$) in the case when $W=0$ can be identified with the Langevin diffusion, the Wasserstein gradient flow of ($\mathsf{pE}$) in the case when $W\neq 0$ corresponds to a (pairwise) McKean–Vlasov SDE, i.e. an SDE whose coefficients depend on the marginal law of the process, given below as
@@
where $\{B^{i}:i\in[N]\}$ is a collection of independent Brownian motions. Moreover, the error from approximating the mean-field limit via this finite particle system has been studied in the literature on propagation of chaos [sznitman1991topics]. Similarly, the Wasserstein gradient flow for ($\mathsf{gE}$) corresponds to the mean-field Langevin dynamics and admits an analogous particle approximation.
The bounds for propagation of chaos have been refined over time, with [lacker2023sharp] recently establishing a tight error dependence $\mathcal{O}(\nicefrac{{1}}{{N}})$ on the total number of particles $N$. These bounds, however, do not translate immediately into algorithmic guarantees. Existing sampling analyses study the propagation of chaos and discretization as a single entangled problem, which thus far have only been able to use weaker ${\mathcal{O}}(\sqrt{\nicefrac{{1}}{{N}}})$ rates for the former. Furthermore, there has been recent interest in using more sophisticated particle-based algorithms, e.g., “non-linear” Hamiltonian Monte Carlo [bou2023nonlinear] and the mean-field underdamped Langevin dynamics [fu2023mean] to reduce the discretization error. Currently, this requires repeatedly carrying out the propagation of chaos and time discretization analyses from the ground up for each instance.
This motivates us to pose the following questions: (1) Can we incorporate improvements in the propagation of chaos literature, such as the $\mathcal{O}(\nicefrac{{1}}{{N}})$ error dependence shown in [lacker2023sharp], to improve existing theoretical guarantees? (2) Can we leverage recent advances in the theory of log-concave sampling to design better algorithms?
Our main proposal in this work is to decouple the error into two terms, representing the propagation of chaos and discretization errors respectively. This simple and modular approach immediately allows us to answer both questions in the affirmative. Namely, we show how to combine established propagation of chaos bounds in various settings [[, including the sharp rate of]]lacker2023sharp with a large class of sophisticated off-the-shelf log-concave samplers, such as interacting versions of the randomized midpoint discretization of the underdamped Langevin dynamics [shen2019randomized, he2020ergodicity], Metropolis-adjusted algorithms [chewi2021optimal, wu2022minimax, altschuler2023faster], and the proximal sampler [LeeSheTia21RGO, chen22improved, fan2023improved]. Our framework yields improvements upon prior state-of-the-art, such as [bou2023nonlinear, fu2023mean], and provides a clear path for future ones.
@@ 1.1 Contributions and Organization
We provide three propagation of chaos results which hold in the $\mathcal{W}_{2}$, $\sqrt{\operatorname{\mathsf{KL}}}$, and $\sqrt{\operatorname{\mathsf{FI}}}$ “metrics”; the rates reflect the distance of the $k$-particle marginal of the finite-particle system from $\pi^{\otimes k}$: (1) In the setting of ($\mathsf{pE}$), under strong displacement convexity, we obtain a $\mathcal{O}(\sqrt{\nicefrac{{k}}{{N}}})$ rate by adapting techniques from [sznitman1991topics, malrieu2001logarithmic]; (2) without assuming displacement convexity, but assuming a weaker interaction, we obtain the sharp rate of $\widetilde{\mathcal{O}}(\nicefrac{{k}}{{N}})$ following [lacker2023sharp]; (3) finally, in the general setting of ($\mathsf{gE}$), and assuming $\mathcal{F}$ is convex along linear interpolations, we obtain a $\mathcal{O}(\sqrt{\nicefrac{{k}}{{N}}})$ rate following [chen2022uniform].
@@ 1.2 Related Work
The McKean–Vlasov SDE was first formulated in the works [mckean1966class, funaki1984certain, Mel1996Interacting], with origins dating to much earlier [boltzmann1872more]. It has applications in many domains, from fluid dynamics [villani2002review] to game theory [lasry2007mean, carmona2018probabilistic]; see [chaintron2022propagationpartone, chaintron2022propagationparttwo] for a comprehensive survey. The kinetic version of this equation is known as the Boltzmann equation, and propagation of chaos has similarly been studied under a variety of assumptions [bolley2010trend, monmarche2017long, guillin2021uniform, guillin2022convergence]. One prominent application within machine learning is the study of infinitely wide two-layer neural networks in the mean-field regime (see §4.2).
@@
The original propagation of chaos arguments of [sznitman1991topics] were first made uniform in time in [malrieu2001logarithmic, malrieu2003convergence] in both entropy and $\mathcal{W}_{2}$. The aforementioned works all achieve an error of order $\widetilde{\mathcal{O}}(\sqrt{\nicefrac{{k}}{{N}}})$, and require a strong convexity assumption on $V$ and $W$. These were later adapted for non-smooth potentials [jabin2017mean, jabin2018quantitative, bresch2023mean]. Finally, [chen2022uniform] obtained an entropic propagation of chaos bound under a higher-order smoothness assumption. See [chaintron2022propagationpartone] for a more complete bibliography.
The breakthrough result of [lacker2023hierarchies] obtained the sharp bound of $\widetilde{\mathcal{O}}(\nicefrac{{k}}{{N}})$ when the interaction is sufficiently weak, and this bound was made uniform in time in [lacker2023sharp]. Their approach differs significantly from previous proofs by considering a local analysis based on the recursive BBGKY hierarchy. These results have been extended to other divergences, e.g., the $\chi^{2}$ divergence, but without a uniform-in-time guarantee [hess2023higher]. In addition, [monmarche2024time] showed an extension of this result under a “convexity at infinity” assumption.
The question of sampling from minimizers of ($\mathsf{pE}$) was first studied in [talay1996probabilistic, bossy1997stochastic, antonelli2002rate]. These works focused on the Euler–Maruyama discretization of the finite-particle system ($\mathsf{pMV}_{N}$), under $L^{\infty}$-boundedness of the gradients. Subsequently, the convergence of the Euler–Maruyama scheme has been studied in many works, including but not limited to [bao2022approximations, dos2022simulation, li2023strong]. The strategy of disentangling finite particle error from time discretization also appears in [karimi2024stochastic], which approaches the problem from the perspective of stochastic approximation. This work, however, is not focused on obtaining quantitative guarantees. Finally, [bou2023nonlinear] considered a non-linear version of Hamiltonian Monte Carlo; we give a detailed comparison with their work in §4.
@@
The mean-field (underdamped) Langevin algorithm for minimizing ($\mathsf{gE}$) was proposed and studied in [chen2022uniform, Chen+24PoCKinetic]. Under alternative assumptions (see §3.1.2), they established propagation of chaos with a $\mathcal{O}(\sqrt{\nicefrac{{k}}{{N}}})$ rate, for both the overdamped and the underdamped finite-particle approximations. Recent works from the machine learning community [nitanda2022convex, suzuki2022uniform, fu2023mean, suzuki2023convergence] studied the application of these algorithms for optimizing two-layer neural networks and obtained sampling guarantees. We provide a detailed comparison with their works in §4.2.
@@ 2 Background and Notation
When $\log(1/\pi)$ is $\alpha$-uniformly convex for $\alpha>0$, it follows from the Bakry–Émery condition that $\pi$ satisfies (LSI) with constant $C_{\mathsf{LSI}}\leq 1/\alpha$ [bakry2014analysis, Proposition 5.7.1].
@@
Lastly, we recall that the celebrated Otto calculus interprets the space $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, equipped with the $\mathcal{W}_{2}$ metric, as a formal Riemannian manifold [otto2001geometry]. In particular, the Wasserstein gradient of a functional $\mathcal{L}:\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})\to\mathbb{R}\cup\{\infty\}$ is given as $\nabla_{\mathcal{W}_{2}}\mathcal{L}=\nabla\delta\mathcal{L}$. Here, $\delta\mathcal{L}$ is the first variation defined as follows: for all $\nu_{0},\nu_{1}\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, $\delta\mathcal{L}(\nu_{0}):\mathbb{R}^{d}\to\mathbb{R}$ satisfies
@@
The first variation is defined up to an additive constant, but the Wasserstein gradient is unambiguous. See [AGS] for a rigorous development. As a shorthand, we will write $\delta\mathcal{L}(\nu_{0},x)\coloneqq\delta\mathcal{L}(\nu_{0})(x)$ and similarly $\nabla_{\mathcal{W}_{2}}\mathcal{L}(\nu_{0},x)\coloneqq\nabla_{\mathcal{W}_{2}}\mathcal{L}(\nu_{0})(x)$.
@@ 2.1.1 The Pairwise McKean–Vlasov Setting
In the formalism introduced in the previous section, we note that ($\mathsf{pMV}$) can be interpreted as Wasserstein gradient flow for ($\mathsf{pE}$). In this paper, we refer to ($\mathsf{pMV}$) as the pairwise McKean–Vlasov process. As noted in the introduction, it has the stationary distribution (1.1) which minimizes ($\mathsf{pE}$).
Recall also that the equation ($\mathsf{pMV}$) is the mean-field limit of the finite-particle system ($\mathsf{pMV}_{N}$). This $N$-particle system has the following stationary distribution: for $x^{1:N}=[x^{1},\dotsc,x^{N}]\in\mathbb{R}^{d\times N}$,
@@ 2.1.2 The General McKean–Vlasov Setting
More generally, we consider the functional ($\mathsf{gE}$) where $\mathcal{F}$ is of the form $\mathcal{F}(\mu)=\mathcal{F}_{0}(\mu)+\frac{\lambda}{2}\int\lVert\cdot\rVert^{2}\,\mathrm{d}\mu$ with $\lambda\geq 0$. The second term acts as regularization and is common in the literature [fu2023mean, suzuki2023convergence]. We can describe its Wasserstein gradient flow as the marginal law of a particle trajectory satisfying the following SDE, which we call the general McKean–Vlasov equation:
@@
and $\rho_{x^{1:N}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{i}}$ is the empirical measure of the particle system. The stationary distribution for ($\mathsf{gMV}_{N}$) is given as follows [chen2022uniform, (2.16)]: for $x^{1:N}=[x^{1},\dotsc,x^{N}]\in\mathbb{R}^{d\times N}$,